# DACAD: Domain Adaptation Contrastive Learning for Anomaly Detection in Multivariate Time Series

### v2

@@ Abstract
In time series anomaly detection (TSAD), the scarcity of labeled data poses a challenge to the development of accurate models. Unsupervised domain adaptation (UDA) offers a solution by leveraging labeled data from a related domain to detect anomalies in an unlabeled target domain. However, existing UDA methods assume consistent anomalous classes across domains. To address this limitation, we propose a novel Domain Adaptation Contrastive learning model for Anomaly Detection in multivariate time series (DACAD), combining UDA with contrastive learning. DACAD utilizes an anomaly injection mechanism that enhances generalization across unseen anomalous classes, improving adaptability and robustness. Additionally, our model employs supervised contrastive loss for the source domain and self-supervised contrastive triplet loss for the target domain, ensuring comprehensive feature representation learning and domain-invariant feature extraction. Finally, an effective Centre-based Entropy Classifier (CEC) accurately learns normal boundaries in the source domain. Extensive evaluations on multiple real-world datasets and a synthetic dataset highlight DACAD’s superior performance in transferring knowledge across domains and mitigating the challenge of limited labeled data in TSAD.
@@ I Introduction
Unsupervised Domain Adaptation (UDA) is a technique used to transfer knowledge from a labelled source domain to an unlabeled target domain. Such an approach is particularly useful when labelled data in the target domain are scarce or unavailable [1]. Deep learning methods have become the predominant approach in UDA, offering advanced capabilities and significantly improved performance compared to traditional techniques. UDA is crucial in situations where the performance of deep models drops significantly due to the discrepancy between the data distributions in the source and target domains, a phenomenon known as domain shift [2].
Applications of UDA are diverse, ranging from image and video analysis to natural language processing and time series data analysis. However, the time series domain differs greatly from the image and text data domains for which UDA methods are well developed [3, 4]. UDA can be particularly challenging for time series analysis, given that often i) the aspects of time series that may be relevant to different tasks can vary greatly, such as frequency, magnitude, rate of change in either of these, global shape or local shape [5, 2, 6], and ii) the number of anomalous classes changes between the source and target domains.
@@
Furthermore, in the realm of time series anomaly detection, in the recent model called ContextDA [13], the discriminator aligns source/target domain windows without leveraging label information in the source domain, which makes the alignment ineffective. Our model leverages source labels and anomaly injection for better feature extraction, enhancing the alignment of normal samples. This is particularly vital due to the challenges in aligning anomalies across domains with different anomalous classes.
In our study, we introduce Domain Adaptation Contrastive learning model for Anomaly Detection in time series (DACAD), a unique framework for UDA in multivariate time series anomaly detection, which leverages contrastive learning (CL) and focuses on contextual representations. It forms positive pairs based on proximity and negative pairs using anomaly injection, subsequently learning their embeddings with both supervised CL in the source domain and self-supervised CL in the target domain. To sum up, we make the following contributions:
  • We introduce DACAD, a pioneering CL framework for multivariate time series anomaly detection with UDA. It is a novel deep UDA model based on contrastive representation utilizing a labeled source dataset and synthetic anomaly injection to overcome the lack of labeled data in TSAD, distinguishing between normal and anomalous patterns effectively.
  • Extending on the deep one-class classifier (DeepSVDD) [14], our proposed one-class classifier, the CEC, operates on the principle of spatial separation in the feature space by leveraging the existing label information in the source domain, aiming to bring “normal” sample representations closer to the center and distancing anomalous ones. This spatial adjustment is measured using a distance metric, establishing the basis for anomaly detection.
  • Our comprehensive evaluation with real-world datasets highlights DACAD’s efficiency. In comparison with the recent TSAD deep models and UDA models for time series classification and UDA for anomaly detection, DACAD demonstrates superior performance, emphasizing the considerable importance of our study.
@@ II Related Works
Time series specific UDA approaches like variational recurrent adversarial deep domain adaptation (VRADA) [7], pioneered UDA for multivariate time series, utilizing adversarial learning with an LSTM networks [23] and variational RNN [24] feature extractor. Convolutional deep domain adaptation for time series (CoDATS) [10], built on VRADA’s adversarial training but employed a convolutional neural network as the feature extractor. A metric-based method, time series sparse associative structure alignment (TS-SASA) [8] aligns intra and inter-variable attention mechanisms between domains using MMD. Adding to these, the CL for UDA of time series (CLUDA) [9] model offers a novel approach, enhancing domain adaptation capabilities in time series. All these methods share a common objective of aligning features across domains, each contributing unique strategies to the complex challenge of domain adaptation in time series classification data. However, they are ineffective when applied to TSAD tasks. Additionally, ContextDA [13], is a TSAD model that applies deep reinforcement learning to optimize domain adaptation, framing context sampling as a Markov decision process. However, it is ineffective when the anomaly classes change between the source and the target.
@@
Recently, Large Language Models (LLMs) and Pretrained Models (PMs) have gained attention for addressing Unsupervised Domain Adaptation (UDA) in time series tasks [25, 26]. These models leverage their robust transfer learning capabilities and the benefits of extensive pre-training on a vast array of sequential data across different domains. This pre-training phase equips the models with comprehensive knowledge and advanced pattern recognition skills, allowing them to effectively adapt to and perform well on specific time series tasks through a relatively straightforward fine-tuning process [27, 28]. This extensive pre-training enables LLMs and PMs to adeptly handle time series anomaly detection, even when these tasks differ significantly from the domains on which they were originally trained. Additionally, diffusion model-based UDA methods have emerged as a promising approach, utilizing the inherent strengths of diffusion models in capturing dynamic changes over time [29]. By integrating a small amount of target domain data to guide the adaptation process, these models can achieve effective domain transfer [30, 31]. However, both LLMs and diffusion models are not without their challenges. Issues such as computational efficiency and training stability are prevalent and continue to be significant areas of research and discussion in the field [29, 28].
@@
Deep Learning for TSAD: The field of Time Series Anomaly Detection (TSAD) has advanced significantly, embracing a variety of methods ranging from basic statistical approaches to sophisticated deep learning techniques [32, 33, 34]. Notably, deep learning has emerged as a promising approach due to its autonomous feature extraction capabilities [35]. TSAD primarily focuses on unsupervised [36, 37, 38, 39] and semi-supervised methods [40, 41, 42, 43] to tackle the challenge of limited labelled data availability. Unsupervised methods like OmniAnomaly [44], GDN [45], and BeatGAN [46] are especially valuable in scenarios with sparse anomaly labels, whereas semi-supervised methods leverage available labels effectively.
@@
Advanced models such as LSTM-NDT [37] and THOC [47] excel in minimizing forecasting errors and capturing temporal dependencies. However, models like USAD [38] face challenges with long time series data due to error accumulation in decoding. Additionally, unsupervised representation learning, exemplified by SPIRAL [48] and TST [49], shows impressive performance, albeit with scalability issues in long time series. Newer models like TNC [50] and T-Loss [51] aim to overcome these challenges using methods such as time-based negative sampling.
@@
In recent years, several works have been proposed that design universal models for various time series tasks, including anomaly detection. For example, TimesNet [55] transforms 1D time series data into 2D tensors to better capture and model complex temporal variations through multi-periodicity, and it improves performance across various time series analysis tasks. Additionally, some methods based on large models or pre-trained models have been proposed. For example, Zhou et al. [25] leverage pre-trained models, preserving their original self-attention and feedforward layers, to achieve state-of-the-art performance in various time series domain tasks by fine-tuning, demonstrating their effectiveness and universality. There are also specific methods for TSAD using large models or pre-trained models, such as AnomalyLLM [56] and AnomalyBERT [57].
@@ III DACAD
DACAD’s core is CL as inspired by a recent UDA for time series work [9], which strengthens its ability to handle changes between domains by improving the feature representation learning in source and target domains. As described in subsection III-C, in the target domain, we use a self-supervised contrastive loss [58] by forming triplets to minimize the distance between similar samples and maximize the distance between different samples. Additionally, in the source domain, we leverage label information of both anomaly and normal classes and propose an effective supervised contrastive loss [59] named supervised mean-margin contrastive loss. A Temporal Convolutional Network (TCN) in DACAD captures temporal dependencies, generating domain-invariant features. A discriminator ensures these features are domain-agnostic, leading to consistent anomaly detection across domains. DACAD’s strength lies in its dual ability to adapt to new data distributions and to distinguish between normal and anomalous patterns effectively. In DACAD, time series data is split into overlapping windows of size $WS$ with a stride of 1, forming detailed source ($S$) and target ($T$) datasets. Windows in the source domain are classified based on anomaly presence: those with anomalies are marked as anomalous windows ($S_{\text{anom}}$), and others as normal ($S_{\text{norm}}$). Fig. 1 shows the DACAD architecture, and Algorithm 1 details its steps. The following subsections explore DACAD’s components and their functionalities.
Figure 1: DACAD Model Overview: Involves source ($S$) and target ($T$) domains. Source domain uses normal ($S_{\text{norm}}$) and anomalous ($S_{\text{anom}}$) samples, plus synthetic anomalies ($S_{\text{inj}}$) for source triplets ($S_{\text{Triplets}}$) and contrastive loss. Target domain similarly uses proximity-based pair selection and anomaly injection ($T_{\text{inj}}$) to create target triplets ($T_{\text{Triplets}}$). TCN ($\phi^{R}$) is used for feature extraction. Features from both domains are fed into discriminator ($\phi^{D}$) for domain-invariant learning. Source features are classified by classifier $\phi^{CL}$.
@@
12:    Compute $\mathcal{L}_{\text{Cls}}$ using Eq. (III-E) and $\phi^{R}(S)$triangleright (III-E)
@@ III-B Pair Selection
In DACAD’s pair selection step, appropriate triplets from source and target domains are created for CL. In the source domain, we use labels to form distinct lists of normal samples ($S_{\text{norm}}$), anomalous samples ($S_{\text{anom}}$), and anomaly-injected samples ($S_{\text{inj}}$). This allows for a supervised CL approach, enhancing differentiation between normal and anomalous samples. Here, triplets ($S_{\text{triplets}}$) consist of an anchor (normal window from $S_{\text{norm}}$), a positive (different normal window from $S_{\text{norm}}$), and a negative (randomly selected from either an anomalous window from $S_{\text{anom}}$ or an anomaly-injected anchor from $S_{\text{inj}}$).
In the target domain, lacking ground truth labels, a self-supervised approach shapes the triplets ($T_{\text{triplets}}$). Each target triplet includes an anchor (original window from $T$), a positive (temporally close window to anchor, from $T$, likely sharing similar characteristics), and a negative (anomaly-injected anchor from $T_{\text{inj}}$).
@@ III-C Representation Layer ($\phi^{R}$)
In our model, we employ a TCN [61] for the representation layer, which is adept at handling both multivariate time series windows. This choice is motivated by the TCN’s ability to capture temporal dependencies effectively, a critical aspect in time series analysis. The inputs to the TCN are the datasets and triplets from both domains, specifically $S$, $T$, $S_{\text{Triplets}}$, and $T_{\text{Triplets}}$. The outputs of the TCN, representing the transformed feature space, are
@@ III-C2 Self-supervised Contrastive Triplet Loss for Target Domain ($\mathcal{L}_{\text{SelfCont}}$)
For the target domain, we employ a self-supervised contrastive approach using triplet loss [58], designed to ensure that the anchor window is closer to a positive window than to a negative window by a specified margin. The self-supervised triplet loss formula is shown in equation (2):
@@
In this setup, the anchor window $a^{t}_{i}$ is compared to a positive window $p^{t}_{i}$ (a randomly selected window in temporal proximity to the anchor from $T$) and a negative window $n^{t}_{i}$ (the anomaly-injected version of the anchor from $T_{\text{inj}}$), ensuring the anchor is closer to the positive than the negative by at least the margin $m$.
@@ III-D Discriminator Component ($\phi^{D}$)
Designed to differentiate between features extracted from the source and target domains, the discriminator employs adversarial training techniques similar to those found in Generative Adversarial Networks (GANs) [62]. Discriminator $\phi^{D}$ is trained to differentiate between source and target domain features, while $\phi^{R}$ is conditioned to produce domain-invariant features.
@@ III-E Centre-based Entropy Classifier ($\phi^{CL}$)
Extending the DeepSVDD [14] designed for anomaly detection, the CEC ($\phi^{CL}$) in DACAD is proposed as an effective anomaly detection classifier in the source domain. It assigns anomaly scores to time series windows, using labeled data from the source domain $S$ for training and applying the classifier to target domain data $T$ during inference. It is centered around a Multi-Layer Perceptron (MLP) with a unique “centre“ parameter crucial for classification.
The classifier operates by spatially separating transformed time series window representations ($\phi^{R}(w_{i}^{s})$) in the feature space relative to a predefined “centre“ $c$. The MLP aims to draw normal sample representations closer to $c$ and push anomalous ones further away. This spatial reconfiguration is quantified using a distance metric, forming the basis for anomaly scoring. A sample closer to $c$ is considered more normal, and vice versa. The effectiveness of this spatial adjustment is measured using a loss function based on BCE, expressed in equation (III-E):
$\mathcal{L}_{\text{Cls}}=$ $-\frac{1}{|S|}\sum_{i=1}^{|S|}\left[y_{i}\cdot\log(\|g(w_{i}^{s})-c\|_{2}^{2})\right.$
$\left.+(1-y_{i})\cdot\log(1-\|g(w_{i}^{s})-c\|_{2}^{2})\right]$
where $\text{where }g(w)=\phi^{CL}(\phi^{R}(w))$ (4)

### v3

@@ Abstract
In time series anomaly detection (TSAD), the scarcity of labeled data poses a challenge to the development of accurate models. Unsupervised domain adaptation (UDA) offers a solution by leveraging labeled data from a related domain to detect anomalies in an unlabeled target domain. However, existing UDA methods assume consistent anomalous classes across domains. To address this limitation, we propose a novel Domain Adaptation Contrastive learning model for Anomaly Detection in multivariate time series (DACAD), combining UDA with contrastive learning. DACAD utilizes an anomaly injection mechanism that enhances generalization across unseen anomalous classes, improving adaptability and robustness. Additionally, our model employs supervised contrastive loss for the source domain and self-supervised contrastive triplet loss for the target domain, ensuring comprehensive feature representation learning and domain-invariant feature extraction. Finally, an effective Center-based Entropy Classifier (CEC) accurately learns normal boundaries in the source domain. Extensive evaluations on multiple real-world datasets and a synthetic dataset highlight DACAD’s superior performance in transferring knowledge across domains and mitigating the challenge of limited labeled data in TSAD.
@@ I Introduction
Unsupervised Domain Adaptation (UDA) is a technique used to transfer knowledge from a labeled source domain to an unlabeled target domain. Such an approach is particularly useful when labeled data in the target domain is scarce or unavailable [1]. Deep learning methods have become the predominant approach in UDA, offering advanced capabilities and significantly improved performance compared to traditional techniques. UDA is crucial in situations where the performance of deep models drops significantly due to the discrepancy between the data distributions in the source and target domains, a phenomenon known as domain shift [2].
Applications of UDA are diverse, ranging from image and video analysis to natural language processing and time series data analysis. However, the time series domain differs greatly from the image and text data domains for which UDA methods are well developed [3, 4]. UDA can be particularly challenging for time series analysis, given that often i) the aspects of time series that may be relevant to different tasks can vary greatly, such as frequency, magnitude, rate of change in either of these, global shape or local shape [5, 2, 6] and ii) the number of anomalous classes changes between the source and target domains.
@@
To further substantiate our approach, theoretical insights indicate that reducing contrastive loss lowers the Class-wise Mean Maximum Discrepancy (CMMD) [13], thereby improving domain adaptation. This reinforces the validity of integrating contrastive learning with UDA in our approach and strengthens the theoretical foundation of our work.
Furthermore, in the realm of time series anomaly detection, in the recent model called ContextDA [14], the discriminator aligns source/target domain windows without leveraging label information in the source domain, which makes the alignment ineffective. Unlike these approaches, our proposed model leverages source labels and anomaly injection to enhance feature extraction, particularly by focusing on distinguishing normal from anomalous samples through contrastive learning. This anomaly injection follows the method that has been proposed in the CARLA model [15], demonstrating its efficacy in real-world scenarios. Our model leverages source labels and anomaly injection for better feature extraction, enhancing the alignment of normal samples. This is particularly vital due to the challenges in aligning anomalies across domains with different anomalous classes.
In our study, we introduce the Domain Adaptation Contrastive learning model for Anomaly Detection in time series (DACAD), a unique framework for UDA in multivariate time series (MTS) anomaly detection, which leverages contrastive learning (CL) and focuses on contextual representations. It forms positive pairs based on proximity and negative pairs using anomaly injection, subsequently learning their embeddings with both supervised CL in the source domain and self-supervised CL in the target domain. To sum up, we make the following contributions:
  • We introduce DACAD, a pioneering contrastive learning framework for multivariate time series anomaly detection with UDA. DACAD is a novel deep UDA model that incorporates both supervised contrastive learning in the source domain and self-supervised contrastive learning in the target domain. To address the scarcity of real anomalies, we enhance representation learning through synthetic anomaly injection, providing effective differentiation between normal and anomalous patterns. The source code of DACAD is available on GitHub [^1] https://github.com/zamanzadeh/DACAD.
  • Our proposed Center-based Entropy Classifier (CEC) introduces a novel enhancement to DeepSVDD [16] by explicitly leveraging source domain label information to refine anomaly detection. CEC enforces spatial separation in the feature space by pulling “normal” sample representations closer to the center and distancing anomalous ones. This spatial adjustment is guided by a distance metric that enhances anomaly detection by improving class separability.
  • Our comprehensive evaluation with real-world datasets highlights DACAD’s efficiency. In comparison with the recent TSAD deep models and UDA models for time series classification and UDA for anomaly detection, DACAD demonstrates superior performance, emphasizing the effective contribution of our approach.
@@ II Related Works
Time series-specific UDA approaches like variational recurrent adversarial deep domain adaptation (VRADA) [7], pioneered UDA for MTS, utilizing adversarial learning with an LSTM network [25] and variational RNN [26] feature extractor. Convolutional deep domain adaptation for time series (CoDATS) [10], built on VRADA’s adversarial training, but employed a convolutional neural network as the feature extractor. A metric-based method, time series sparse associative structure alignment (TS-SASA) [8] aligns intra and inter-variable attention mechanisms between domains using MMD. Adding to these, the CL for UDA of time series (CLUDA) [9] model offers a novel approach, enhancing domain adaptation capabilities in time series. All these methods share a common objective of aligning features across domains, each contributing unique strategies to the complex challenge of domain adaptation in time series classification data. However, they are ineffective when applied to TSAD tasks. Additionally, ContextDA [14] is a TSAD model that applies deep reinforcement learning to optimize domain adaptation, framing context sampling as a Markov decision process. However, it is ineffective when the anomaly classes change between the source and the target.
@@
Recently, Large Language Models (LLMs) and Pretrained Models (PMs) have gained attention for addressing Unsupervised Domain Adaptation (UDA) in time series tasks [27, 28]. These models leverage their robust transfer learning capabilities and the benefits of extensive pre-training on a vast array of sequential data across different domains. This pre-training phase equips the models with comprehensive knowledge and advanced pattern recognition skills, allowing them to effectively adapt to and perform well on specific time series tasks through a relatively straightforward fine-tuning process [29, 30]. This extensive pre-training enables LLMs and PMs to adeptly handle time series anomaly detection, even when these tasks differ significantly from the domains on which they were originally trained. Additionally, diffusion model-based UDA methods have emerged as a promising approach, utilizing the inherent strengths of diffusion models in capturing dynamic changes over time [31]. By integrating a small amount of target domain data to guide the adaptation process, these models can achieve effective domain transfer [32, 33]. However, neither LLM nor diffusion models are without their challenges. Issues such as computational efficiency and training stability are prevalent and continue to be significant areas of research and discussion in the field [31, 30].
@@
Deep Learning for TSAD: The field of Time Series Anomaly Detection (TSAD) has advanced significantly, embracing a variety of methods ranging from basic statistical approaches to sophisticated deep learning techniques [34, 35, 36]. Notably, deep learning has emerged as a promising approach due to its autonomous feature extraction capabilities [37]. TSAD primarily focuses on unsupervised [38, 39, 40, 41] and semi-supervised methods [42, 43, 44] to tackle the challenge of limited labeled data availability. Unsupervised methods like OmniAnomaly [45] and GDN [46] are especially valuable in scenarios with sparse anomaly labels, whereas semi-supervised methods leverage available labels effectively.
@@
Advanced models such as LSTM-NDT [39] and THOC [47] excel in minimizing forecasting errors and capturing temporal dependencies. However, models like USAD [40] face challenges with long time series data due to error accumulation in decoding. Additionally, unsupervised representation learning, exemplified by SPIRAL [48] and TST [49], shows impressive performance, albeit with scalability issues in long time series. Newer models like TNC [50] aim to overcome these challenges using methods such as time-based negative sampling.
@@ III DACAD
DACAD’s core is CL as inspired by a recent UDA for time series work [9], which strengthens its ability to handle changes between domains by improving the feature representation learning in source and target domains. As described in subsection III-C, in the target domain, we use a self-supervised contrastive loss [54] by forming triplets to minimize the distance between similar samples and maximize the distance between different samples. Additionally, in the source domain, we leverage label information of both anomaly and normal classes and propose an effective supervised contrastive loss [55] named supervised mean-margin contrastive loss. A Temporal Convolutional Network (TCN) in DACAD captures temporal dependencies, generating domain-invariant features. A discriminator ensures these features are domain-agnostic, leading to consistent anomaly detection across domains. DACAD’s strength lies in its dual ability to adapt to new data distributions and to distinguish between normal and anomalous patterns effectively. In DACAD, time series data is split into overlapping windows of size $W_{S}$ with a stride of 1, forming detailed source ($S$) and target ($T$) datasets. Source windows are classified as normal ($S_{\text{norm}}$) anomalous $S_{\text{anom}}$) based on anomaly presence. Fig. 1 shows the DACAD architecture, and Algorithm 1 details its steps. The following subsections explore DACAD’s components and their functionalities.
Figure 1: DACAD Model Overview: Involves source ($S$) and target ($T$) domains. Source domain uses normal ($S_{\text{norm}}$) and anomalous ($S_{\text{anom}}$) samples, plus synthetic anomalies ($S_{\text{inj}}$) for source triplets ($S_{\text{Triplets}}$) and contrastive loss. Target domain similarly uses proximity-based pair selection and anomaly injection ($T_{\text{inj}}$) to create target triplets ($T_{\text{Triplets}}$). TCN ($\phi^{R}$) is used for feature extraction. Features from both domains are fed into the discriminator ($\phi^{D}$) for domain-invariant learning. Source features are classified by classifier $\phi^{CL}$.
@@
12:    Compute $\mathcal{L}_{\text{Cls}}$ using Eq. (4) and $\phi^{R}(S)$triangleright (III-E)
@@ III-B Pair Selection
In DACAD’s pair selection step, appropriate triplets from the source and target domains are created for CL. In the source domain, we use labels to form distinct lists of normal samples ($S_{\text{norm}}$), anomalous samples ($S_{\text{anom}}$), and anomaly-injected samples ($S_{\text{inj}}$). This allows for a supervised CL approach, enhancing differentiation between normal and anomalous samples. Here, triplets ($S_{\text{triplets}}$) consist of an anchor (normal window from $S_{\text{norm}}$), a positive (different normal window from $S_{\text{norm}}$), and a negative (randomly selected from either an anomalous window from $S_{\text{anom}}$ or an anomaly-injected anchor from $S_{\text{inj}}$).
In the unlabeled target domain, a self-supervised approach constructs the triplets ($T_{\text{triplets}}$). Each target triplet includes an anchor (original window from $T$), a positive (temporally close window to anchor, from $T$, likely sharing similar characteristics), and a negative (anomaly-injected anchor from $T_{\text{inj}}$).
@@ III-C Representation Layer ($\phi^{R}$)
In our model, we employ a TCN [56] for the representation layer, which is adept at handling MTS windows. This choice is motivated by the TCN’s ability to capture temporal dependencies effectively, a critical aspect in time series analysis. The inputs to the TCN are the datasets and triplets from both domains, specifically $S$, $T$, $S_{\text{Triplets}}$, and $T_{\text{Triplets}}$. The outputs of the TCN, representing the transformed feature space, are
@@ III-C2 Self-supervised Contrastive Triplet Loss for Target Domain ($\mathcal{L}_{\text{SelfCont}}$)
For the target domain, we employ a self-supervised contrastive approach using triplet loss [54], designed to ensure that the anchor window is closer to a positive window than to a negative window by a specified margin. The anchor is assumed to be normal, following the standard TSAD assumption that the majority of the data is normal, as anomalies are rare by nature. The self-supervised triplet loss formula is shown in Equation (2):
@@
In this setup, the anchor window $a^{t}_{i}$ is compared to a positive window $p^{t}_{i}$ (a randomly selected nearby window from $T$) and a negative window $n^{t}_{i}$ (the anomaly-injected version of the anchor from $T_{\text{inj}}$), ensuring the anchor is closer to the positive than the negative by at least the margin $m$.
@@ III-D Discriminator Component ($\phi^{D}$)
Designed to differentiate between features extracted from the source and target domains, the discriminator employs adversarial training techniques similar to those found in Generative Adversarial Networks [57]. Discriminator $\phi^{D}$ is trained to differentiate between source and target domain features, while $\phi^{R}$ is conditioned to produce domain-invariant features.
@@ III-E Centre-based Entropy Classifier ($\phi^{CL}$)
Extending the DeepSVDD [16] designed for anomaly detection, the CEC ($\phi^{CL}$) in DACAD is proposed as an effective anomaly detection classifier in the source domain. It assigns anomaly scores to time series windows, using labeled data from the source domain $S$ for training and applying the classifier to target domain data $T$ during inference. It is centered around a Multi-Layer Perceptron (MLP) with a unique “center” parameter crucial for classification.
The classifier operates by spatially separating transformed time series window representations ($\phi^{R}(w_{i}^{s})$) in the feature space relative to a learnable “center” $c$. The MLP aims to draw normal sample representations closer to $c$ and push anomalous ones further away. This spatial reconfiguration is quantified using a distance metric, forming the basis for anomaly scoring. A sample closer to $c$ is considered more normal, and vice versa. The effectiveness of this spatial adjustment is measured using a BCE-based loss function, expressed in Equation (4):
$\mathcal{L}_{\text{Cls}}=$ (4)
$-\frac{1}{|S|}\sum_{i=1}^{|S|}\left[y_{i}\cdot\log(\|g(w_{i}^{s})-c\|_{2}^{2})+(1-y_{i})\cdot\log(1-\|g(w_{i}^{s})-c\|_{2}^{2})\right]$
$where~{}g(w)=\phi^{CL}(\phi^{R}(w))$